OpenStudy (anonymous):

7/33 x y=-56/3 what dose y=?

OpenStudy (anonymous):

u mean does?

jhonyy9 (jhonyy9):

7/33 xy= -56/3 7xy= -56*11 y= -56*11/7x = - 8*7*11/7x = - 88/x